# A new twist on the Majorana surface code: Bosonic and fermionic defects for fault-tolerant quantum computation

### v3

@@ I Introduction
The rest of this work is structured as follows. We start, in Section II, with the main ingredients of MSCs, including their anyons, and fault-tolerant quantum computation. We also establish here the types of boundaries supported by MSCs. In Section III, we classify MSC twists and detail their properties, including their logical Hilbert space dimensions (i.e., their “quantum dimensions”). As in bosonic codes, twists are classified via the symmetries of the MSC’s anyons. But logical MZMs are inherently fermionic and hence have no counterparts in bosonic codes. We find that a phase transition along the domain wall connecting two twists can result in logical MZMs emerging at the twists. This expands on how one-dimensional physics can enrich the classification of twist defects [Defects_Abelian_states], by showing how this may lead to non-local logical MZMs. While mentioned in the MSC literature [Maj_Ferm_Codes, Maj_Triangle_Code2018], the range of scenarios in which one expects to find logical MZMs and their link to twists have not yet been explored.
In Section IV, we discuss how universal fault-tolerant quantum computation can be performed with twists in the MSC, simultaneously utilizing both logical qubits and logical MZMs. This allows for roughly twice the information to be stored and manipulated per twist as would be possible in bosonic approaches. The twists we discuss provide independent ways to protect information from fermion-parity-violating errors (so-called quasi-particle poisoning or QPP events [MZM_TQC_Review2015]) and fermion-parity-conserving (FPC) errors.
In Section V we explore the spatial overheads of computing with logical MZMs, and show that these may be considerably reduced if the QPP probability is suitably small. We focus on a fault-tolerant implementation of “Pauli-based computation” (PBC) [Bravyi_PBC2016, Mithuna_Magic_State_PBC2019, FPBC] with the MSC; PBC is a measurement-based model of quantum computation that optimizes the use of quantum resources, at the cost of additional, but efficient, classical computing.
Finally, in Section VI, we present a new method to implement fault-tolerant gates in MSCs. This is a fermionic analogue of transversal gates in bosonic codes. It implements encoded Clifford gates by braiding the MSC’s constituent MZMs. This technique is particularly suitable for small codes, and hence could be utilized in early-generation Majorana devices. In Section LABEL:sec:Conclusion we conclude and discuss avenues for future work.
II Basic MSC ingredients
@@
II.1 Majorana Fermion Codes
@@
The code’s logical operators are even-weight elements of Maj$(2n)$ that commute with all $s\in\mathcal{S}$ but are not themselves in $\mathcal{S}$. We define logical operators $\bar{X}_{j}$ and $\bar{Z}_{j}$ such that they act as Pauli operators on logical qubit $j$. Note that a logical operator $L$ and $Ls$ act equivalently on code states for any $s\in\mathcal{S}$. Hence we define the equivalence class of logical operator $L$ to be $[L]=\{Ls|s\in\mathcal{S}\}$. Some Majorana fermion codes also have parity-non-preserving (odd-weight) operators that commute with all members of $\mathcal{S}$ apart from $\Gamma$ (in such cases $\Gamma$ is removed from $\mathcal{S}$, and we separately require that logical operators commute with $\Gamma$). We refer to these operators as logical MZMs, for reasons mentioned above (see also Section III).
Suppose the code is defined on a $D$-dimensional lattice with one Majorana operator per site. Suppose also that $\mathcal{S}$ has generators $\mathcal{O}_{p}$ whose support (i.e., sites where $\mathcal{O}_{p}$ differs from identity) is local. While logical operators in $[L]$ act the same way on code states, they can have different weights. The distance of the code is defined as the smallest among the weights of logical operators, $\min_{L,s}|Ls|$. To further characterize codes, we define the diameter $\text{diam}(\Gamma_{A})$ of a logical operator $\Gamma_{A}$ to be the geometrical diameter of its support. Similarly to the distance, the code diameter is defined as $\min_{L,s}\text{diam}(Ls)$. As noted in Ref. Maj_Ferm_Codes, and as we explain in Section III.3, a code with logical MZMs can have large diameter but small distance. See Table 1 for a summary of the objects and quantities introduced above.
@@
II.2 Universal Fault-Tolerant Quantum Computation
@@
There are many proposed methods for performing logic gates fault-tolerantly on information stored in quantum codes. For MSCs, previous proposals have suggested replacing Clifford gates with measurements, which can be performed fault-tolerantly via “lattice surgery” [QC_with_MFCs, Surface_Code_Lattice_Surg, Maj_Triangle_Code2018]. These involve fault-tolerantly measuring extra stabilizer operators, which results in the fusion of two or more patches of code. In the bosonic surface code, Clifford gates can be performed via the introduction of holes or twist defects (see Section III) in the lattice. Using “code deformation”, one can braid these lattice features, thereby achieving gates on the logical qubits [Bombin_Code_Deformation, Bombin_twists_code_deformation, Holes_Twists_SurfaceCode]. Holes can also be introduced to the MSC and braided to achieve CNOT gates [MjFerm_Surface_Code]. The Hadamard gate can also be applied to these qubits, but this involves measuring a high-weight Majorana operator, which may be challenging in physical setups [MjFerm_Surface_Code, Roadmap_to_MSCs]. The $S$ gate can be implemented by preparing an ancilla logical qubit in a high-fidelity Pauli-$Y$ eigenstate, and then performing $H$ and CNOT gates to the ancilla and target qubits [MjFerm_Surface_Code, Roadmap_to_MSCs]. Thus, all Clifford generators can be enacted in MSCs with holes.
@@
II.3 Majorana Surface Code
@@ II.4 Anyons in the MSC
Similarly to anyons, string operators also obey fusion rules. Viewing $\mathsf{C}$ as the endpoint of string operator of color $\mathsf{C}\in\{\mathsf{R},\mathsf{G},\mathsf{B}\}$ and $\mathbf{1}$ as the identity operator, we can apply Equations (II.4) to string operators even away from their endpoints. (We equate $\mathsf{RGB}$ with $\mathbf{1}$ in this context because an $\mathsf{RGB}$ pair can be created by a Majorana pair without string operators.) Fusion and splitting of string operators is implemented by multiplication with stabilizers. For example, the green string operator $\bar{Z}$ in Figure 1 can be split, by multiplication with green stabilizers, into blue and red strings (i.e., products of blue and red bond operators) along its length. This is shown in Figure 3. Hence the non-trivial rules from Equations (II.4) for string operators become $\mathsf{R}=\mathsf{G}\times\mathsf{B}$ and all cyclic permutations of this.
II.5 Boundaries in the MSC
@@
III Twist Defects of the Majorana Surface Code
@@
Twists in the MSC differ from those in bosonic models in an important way: they can be further subdivided into “bosonic” and “fermionic” types of twist. These differ by a single MZM. However, rather than an extra MZM simply furnishing the twists, as might be expected, in Figure 5 we show that in fact the result can be a logical MZM located at the twist (cf. Section II.1). In Figure 5a we provide a lattice realization of bosonic $Z_{2}$ and $Z_{3}$ twists and the domain walls connecting them, while in Figure 5b we show $Z_{2}$ and $Z_{3}$ fermionic twists. In the latter, the lattice contains pentagonal holes. The (Hermitian) product of all Majorana operators around the boundary of this hole is a fermion-parity-odd operator that, by Equation (2), commutes with all plaquette operators. These operators cannot be included in the stabilizer group, owing to their odd fermion parity. Instead they are logical MZMs. They generalize the key features of MZMs, including their commutation with the Hamiltonian. We discuss these logical MZMs appearing at fermionic twists below, in Section III.3.
@@
When the MSC is defined on the honeycomb lattice, some Majoranas along domain walls are not included in any plaquettes. In Figure 5a and 5b, we draw orange, pill-shape faces between pairs of Majoranas left out of the lattice in this way. These Majorana operators furnish extraneous fermion modes (cf. Section II.1), so we introduce additional bilinear stabilizers ($i\gamma_{a}\gamma_{b}$, for left-out operators $\gamma_{a}$, $\gamma_{b}$), represented by the orange faces in Figure 5, to remove the ground-state degeneracy associated with these modes, i.e., to remove these states from our code space. Those Majorana operators included in orange stabilizers are not included in any other stabilizer of the code.
@@
III.1 Quantum Dimensions of Twists
@@
III.2 Logical Qubits from Twists
@@
In Figure 7b we illustrate the logical operators for four $Z_{3}$ twists. In the figure, an anyon crossing a domain wall in the direction of the arrows is transformed according to the permutation $T_{\mathsf{RGB}}$, while those crossing in the opposite direction are transformed according to the inverse permutation. As can be seen, four $Z_{3}$ twists permit twice the number of logical operators as four $Z_{2}$ twists, owing to the extra degeneracy they afford (cf. Section III.1).
There is an arbitrariness to the logical operators shown. In Figure 7a, one may wrap blue string operators around the two upper-most or left-most twists (these operators are stabilizer elements). The blue strings can be fused with the logical operators indicated, in the sense outlined in Section II.4, thereby changing the red string operators to green ones and vice versa. If the code is embedded on a sphere, red and green string operators encircling all four twists are stabilizers, since they may be deformed around to the other side of the sphere, where they are not encircling any twists. With different boundary conditions, this becomes complicated, and we deal with these cases in the following section. For now, we continue with the simple case of a sphere. Operators $\bar{X}$ and $\bar{Z}$ can then be deformed into operators that instead circle the two bottom and right-most twists respectively. In Figure 7b, one cannot change the colors of the strings, since any string operator encircling a pair of $Z_{3}$ twists is not a stabilizer. But on a sphere, all strings may be deformed to encircle the other two twists in the set of four.
@@
III.3 Logical MZMs from Fermionic Twists
@@
Both bosonic and fermionic twists in the MSC host the BT qubits discussed above. This same type of topological degeneracy is seen in twists of qubit-based codes, which is why we refer to it as “bosonic.” For fermionic twists in the MSC we will label the corresponding bosonic logical operators as $\bar{X}_{B}$ and $\bar{Z}_{B}$. In addition, fermionic twists also host logical MZMs. In the same way as four physical MZMs can store a single qubit in a fermionic parity sector [MZM_TQC_Review2015], so too can four logical MZMs $\Upsilon_{a},\Upsilon_{b},\Upsilon_{c},\Upsilon_{d}$. We can also define encoded Pauli operators acting on the qubits stored by these logical MZMs. For example, for four logical MZMs, we may have “fermionic” logical operators $\bar{X}_{F}=i\Upsilon_{a}\Upsilon_{b}$ and $\bar{Z}_{F}=i\Upsilon_{a}\Upsilon_{c}$. We will refer to qubits stored in this way as “fermionic twist (FT) qubits.” Figure 8a shows the operator $i\Upsilon_{a}\Upsilon_{b}$ for two twists, deformed such that it is the product of a red string operator encircling the twists, and a blue string operator terminating on the two twists. It also shows the representation in terms of string operators of a single logical Majorana $\Upsilon_{c}$, which is equivalent to the localization of an $\mathsf{RGB}$ anyon at that twist (cf. Section III.1).
The logical operators $\bar{X}_{F}$ and $\bar{Z}_{F}$ have low minimum weights but large diameters, if the twists are far-separated (cf. Section II.1). If only local, fermion parity conserving (FPC, i.e., even-weight) errors can occur, FT qubits are protected by the twists’ separation [Maj_Ferm_Codes]. (Here we assume local coupling to the environment so that elementary FPC error operators are locally supported; all possible FPC error operators are products of these elementary FPC operators.) In realistic setups, however, we also have to contend with quasi-particle poisoning (QPP) resulting in non-FPC errors. The simplest such error operator is a single Majorana operator. To protect FT qubits from QPP, we can grow the size of the odd-length boundaries associated with twists and hence grow the weights of logical MZMs. As such (assuming we also increase twist separation), the FT qubits will have larger distances and so an increased protection against QPP noise. An example of such an enlarged hole is shown in Figure 8b.
@@
IV Fault-Tolerant Computation With Twists
In this section we detail how to perform fault-tolerant quantum computation with FT and BT qubits. These procedures will require the ability to move twists via “code deformation” (Section IV.1). In Section IV.2 we discuss how to prepare logical Pauli eigenstates and apply Clifford gates to BT qubits. In Sections IV.3 and IV.5 we discuss performing measurements of BT and FT logical operators. The above, combined with the preparation of magic states (Section IV.4), is sufficient for universal computation with both types of qubit.
IV.1 Code Deformation

### v4

@@ 1 Introduction
The rest of this work is structured as follows. We start, in Section 2, with the main ingredients of MSCs, including their anyons, and fault-tolerant quantum computation. We also establish here the types of boundaries supported by MSCs. In Section 3, we classify MSC twists and detail their properties, including their logical Hilbert space dimensions (i.e., their “quantum dimensions”). As in bosonic codes, twists are classified via the symmetries of the MSC’s anyons. But logical MZMs are inherently fermionic and hence have no counterparts in bosonic codes. We find that a phase transition along the domain wall connecting two twists can result in logical MZMs emerging at the twists. This expands on how one-dimensional physics can enrich the classification of twist defects [Defects_Abelian_states], by showing how this may lead to non-local logical MZMs. While mentioned in the MSC literature [Maj_Ferm_Codes, Maj_Triangle_Code2018], the range of scenarios in which one expects to find logical MZMs and their link to twists have not yet been explored.
In Section 4, we discuss how universal fault-tolerant quantum computation can be performed with twists in the MSC, simultaneously utilizing both logical qubits and logical MZMs. This allows for roughly twice the information to be stored and manipulated per twist as would be possible in bosonic approaches. The twists we discuss provide independent ways to protect information from fermion-parity-violating errors (so-called quasi-particle poisoning or QPP events [MZM_TQC_Review2015]) and fermion-parity-conserving (FPC) errors.
In Section 5 we explore the spatial overheads of computing with logical MZMs, and show that these may be considerably reduced if the QPP probability is suitably small. We focus on a fault-tolerant implementation of “Pauli-based computation” (PBC) [Bravyi_PBC2016, Mithuna_Magic_State_PBC2019, FPBC] with the MSC; PBC is a measurement-based model of quantum computation that optimizes the use of quantum resources, at the cost of additional, but efficient, classical computing.
Finally, in Section 6, we present a new method to implement fault-tolerant gates in MSCs. This is a fermionic analogue of transversal gates in bosonic codes. It implements encoded Clifford gates by braiding the MSC’s constituent MZMs. This technique is particularly suitable for small codes, and hence could be utilized in early-generation Majorana devices. In Section LABEL:sec:Conclusion we conclude and discuss avenues for future work.
2 Basic MSC Ingredients
@@
2.1 Majorana Fermion Codes
@@
The code’s logical operators are even-weight elements of Maj$(2n)$ that commute with all $s\in\mathcal{S}$ but are not themselves in $\mathcal{S}$. We define logical operators $\bar{X}_{j}$ and $\bar{Z}_{j}$ such that they act as Pauli operators on logical qubit $j$. Note that a logical operator $L$ and $Ls$ act equivalently on code states for any $s\in\mathcal{S}$. Hence we define the equivalence class of logical operator $L$ to be $[L]=\{Ls|s\in\mathcal{S}\}$. Some Majorana fermion codes also have parity-non-preserving (odd-weight) operators that commute with all members of $\mathcal{S}$ apart from $\Gamma$ (in such cases $\Gamma$ is removed from $\mathcal{S}$, and we separately require that logical operators commute with $\Gamma$). We refer to these operators as logical MZMs, for reasons mentioned above (see also Section 3).
Suppose the code is defined on a $D$-dimensional lattice with one Majorana operator per site. Suppose also that $\mathcal{S}$ has generators $\mathcal{O}_{p}$ whose support (i.e., sites where $\mathcal{O}_{p}$ differs from identity) is local. While logical operators in $[L]$ act the same way on code states, they can have different weights. The distance of the code is defined as the smallest among the weights of logical operators, $\min_{L,s}|Ls|$. To further characterize codes, we define the diameter $\text{diam}(\Gamma_{A})$ of a logical operator $\Gamma_{A}$ to be the geometrical diameter of its support. Similarly to the distance, the code diameter is defined as $\min_{L,s}\text{diam}(Ls)$. As noted in Ref. [Maj_Ferm_Codes], and as we explain in Section 3.3, a code with logical MZMs can have large diameter but small distance. See Table 1 for a summary of the objects and quantities introduced above.
@@
2.2 Universal Fault-Tolerant Quantum Computation
@@
There are many proposed methods for performing logic gates fault-tolerantly on information stored in quantum codes. For MSCs, previous proposals have suggested replacing Clifford gates with measurements, which can be performed fault-tolerantly via “lattice surgery” [QC_with_MFCs, Surface_Code_Lattice_Surg, Maj_Triangle_Code2018]. These involve fault-tolerantly measuring extra stabilizer operators, which results in the fusion of two or more patches of code. In the bosonic surface code, Clifford gates can be performed via the introduction of holes or twist defects (see Section 3) in the lattice. Using “code deformation”, one can braid these lattice features, thereby achieving gates on the logical qubits [Bombin_Code_Deformation, Bombin_twists_code_deformation, Holes_Twists_SurfaceCode]. Holes can also be introduced to the MSC and braided to achieve CNOT gates [MjFerm_Surface_Code]. The Hadamard gate can also be applied to these qubits, but this involves measuring a high-weight Majorana operator, which may be challenging in physical setups [MjFerm_Surface_Code, Roadmap_to_MSCs]. The $S$ gate can be implemented by preparing an ancilla logical qubit in a high-fidelity Pauli-$Y$ eigenstate, and then performing $H$ and CNOT gates to the ancilla and target qubits [MjFerm_Surface_Code, Roadmap_to_MSCs]. Thus, all Clifford generators can be enacted in MSCs with holes.
@@
2.3 Majorana Surface Code
@@ 2.4 Anyons in the MSC
Similarly to anyons, string operators also obey fusion rules. Viewing $\mathsf{C}$ as the endpoint of string operator of color $\mathsf{C}\in\{\mathsf{R},\mathsf{G},\mathsf{B}\}$ and $\mathbf{1}$ as the identity operator, we can apply Equations (2.4) to string operators even away from their endpoints. (We equate $\mathsf{RGB}$ with $\mathbf{1}$ in this context because an $\mathsf{RGB}$ pair can be created by a Majorana pair without string operators.) Fusion and splitting of string operators is implemented by multiplication with stabilizers. For example, the green string operator $\bar{Z}$ in Figure 1 can be split, by multiplication with green stabilizers, into blue and red strings (i.e., products of blue and red bond operators) along its length. This is shown in Figure 3. Hence the non-trivial rules from Equations (2.4) for string operators become $\mathsf{R}=\mathsf{G}\times\mathsf{B}$ and all cyclic permutations of this.
2.5 Boundaries in the MSC
@@
3 Twist Defects of the Majorana Surface Code
@@
Twists in the MSC differ from those in bosonic models in an important way: they can be further subdivided into “bosonic” and “fermionic” types of twist. These differ by a single MZM. However, rather than an extra MZM simply furnishing the twists, as might be expected, in Figure 5 we show that in fact the result can be a logical MZM located at the twist (cf. Section 2.1). In Figure 5a we provide a lattice realization of bosonic $Z_{2}$ and $Z_{3}$ twists and the domain walls connecting them, while in Figure 5b we show $Z_{2}$ and $Z_{3}$ fermionic twists. In the latter, the lattice contains pentagonal holes. The (Hermitian) product of all Majorana operators around the boundary of this hole is a fermion-parity-odd operator that, by Equation (2), commutes with all plaquette operators. These operators cannot be included in the stabilizer group, owing to their odd fermion parity. Instead they are logical MZMs. They generalize the key features of MZMs, including their commutation with the Hamiltonian. We discuss these logical MZMs appearing at fermionic twists below, in Section 3.3.
@@
When the MSC is defined on the honeycomb lattice, some Majoranas along domain walls are not included in any plaquettes. In Figure 5a and 5b, we draw orange, pill-shape faces between pairs of Majoranas left out of the lattice in this way. These Majorana operators furnish extraneous fermion modes (cf. Section 2.1), so we introduce additional bilinear stabilizers ($i\gamma_{a}\gamma_{b}$, for left-out operators $\gamma_{a}$, $\gamma_{b}$), represented by the orange faces in Figure 5, to remove the ground-state degeneracy associated with these modes, i.e., to remove these states from our code space. Those Majorana operators included in orange stabilizers are not included in any other stabilizer of the code.
@@
3.1 Quantum Dimensions of Twists
@@
3.2 Logical Qubits from Twists
@@
In Figure 7b we illustrate the logical operators for four $Z_{3}$ twists. In the figure, an anyon crossing a domain wall in the direction of the arrows is transformed according to the permutation $T_{\mathsf{RGB}}$, while those crossing in the opposite direction are transformed according to the inverse permutation. As can be seen, four $Z_{3}$ twists permit twice the number of logical operators as four $Z_{2}$ twists, owing to the extra degeneracy they afford (cf. Section 3.1).
There is an arbitrariness to the logical operators shown. In Figure 7a, one may wrap blue string operators around the two upper-most or left-most twists (these operators are stabilizer elements). The blue strings can be fused with the logical operators indicated, in the sense outlined in Section 2.4, thereby changing the red string operators to green ones and vice versa. If the code is embedded on a sphere, red and green string operators encircling all four twists are stabilizers, since they may be deformed around to the other side of the sphere, where they are not encircling any twists. With different boundary conditions, this becomes complicated, and we deal with these cases in the following section. For now, we continue with the simple case of a sphere. Operators $\bar{X}$ and $\bar{Z}$ can then be deformed into operators that instead circle the two bottom and right-most twists respectively. In Figure 7b, one cannot change the colors of the strings, since any string operator encircling a pair of $Z_{3}$ twists is not a stabilizer. But on a sphere, all strings may be deformed to encircle the other two twists in the set of four.
@@
3.3 Logical MZMs from Fermionic Twists
@@
Both bosonic and fermionic twists in the MSC host the BT qubits discussed above. This same type of topological degeneracy is seen in twists of qubit-based codes, which is why we refer to it as “bosonic.” For fermionic twists in the MSC we will label the corresponding bosonic logical operators as $\bar{X}_{B}$ and $\bar{Z}_{B}$. In addition, fermionic twists also host logical MZMs. In the same way as four physical MZMs can store a single qubit in a fermionic parity sector [MZM_TQC_Review2015], so too can four logical MZMs $\Upsilon_{a},\Upsilon_{b},\Upsilon_{c},\Upsilon_{d}$. We can also define encoded Pauli operators acting on the qubits stored by these logical MZMs. For example, for four logical MZMs, we may have “fermionic” logical operators $\bar{X}_{F}=i\Upsilon_{a}\Upsilon_{b}$ and $\bar{Z}_{F}=i\Upsilon_{a}\Upsilon_{c}$. We will refer to qubits stored in this way as “fermionic twist (FT) qubits.” Figure 8a shows the operator $i\Upsilon_{a}\Upsilon_{b}$ for two twists, deformed such that it is the product of a red string operator encircling the twists, and a blue string operator terminating on the two twists. It also shows the representation in terms of string operators of a single logical Majorana $\Upsilon_{c}$, which is equivalent to the localization of an $\mathsf{RGB}$ anyon at that twist (cf. Section 3.1).
The logical operators $\bar{X}_{F}$ and $\bar{Z}_{F}$ have low minimum weights but large diameters, if the twists are far-separated (cf. Section 2.1). If only local, fermion parity conserving (FPC, i.e., even-weight) errors can occur, FT qubits are protected by the twists’ separation [Maj_Ferm_Codes]. (Here we assume local coupling to the environment so that elementary FPC error operators are locally supported; all possible FPC error operators are products of these elementary FPC operators.) In realistic setups, however, we also have to contend with quasi-particle poisoning (QPP) resulting in non-FPC errors. The simplest such error operator is a single Majorana operator. To protect FT qubits from QPP, we can grow the size of the odd-length boundaries associated with twists and hence grow the weights of logical MZMs. As such (assuming we also increase twist separation), the FT qubits will have larger distances and so an increased protection against QPP noise. An example of such an enlarged hole is shown in Figure 8b.
@@
4 Fault-Tolerant Computation With Twists
In this section we detail how to perform fault-tolerant quantum computation with FT and BT qubits. These procedures will require the ability to move twists via “code deformation” (Section 4.1). In Section 4.2 we discuss how to prepare logical Pauli eigenstates and apply Clifford gates to BT qubits. In Sections 4.3 and 4.5 we discuss performing measurements of BT and FT logical operators. The above, combined with the preparation of magic states (Section 4.4), is sufficient for universal computation with both types of qubit.
4.1 Code Deformation